Problem 4: Vehicle Costs (38 points Your aunt hears that youre studying sustainability and energy, and she wants your advice about choosing a fuel-efficient vehicle! She drives 12,000 miles per year (typical for the US) and plans to keep her car for 5 years before selling it (again, quite typical). Shes considering three cars Car Purchase price Tax credit 5-year resale price Fuel efficiency (capital cost)(first year) (residual value) Toyota Yaris $15,000 33 miles/gallon 50 miles/gallon 3 miles/kWh none $10,000 Toyota Prius $24,000 (hybrid) Nissan Leaf $30,000 (electric) $2,500 $12,000 $7,500 $6,000 (f) Fuel usage: For each car, what is the total amount of fuel that she would use per year (in gallons of gasoline or kWh of electricity)? (6 pt) (g) Fuel cost: Assuming that gasoline prices will average $2.00/gallon and Arizona electricity prices will average 13 cents/kWh over the next five years, what would be the annual cost to fuel each of these cars? Which car is the cheapest to drive, based on fuel costs? (3 pt) (h) 5-year cost of ownership: We calculate the lifetime cost of power plants by adding up all the capital and variable costs, and subtracting tax credits and residual value. We can do the same thing for cars to get their cost of ownership. What is the total five-year cost of ownership for these three vehicles, including purchase price, fuel, tax credits, and resale? I didnt give you an exact formula for this but I think you should be able to think it through. (Remember, as we discussed in class, the tax credit is a one-time rebate After you buy an alternative fuel car, the government gives you a break on your taxes So its money that you get back after youve bought the car.) (15 pt) (i) Comparing cars: Which car is the cheapest when you look at the full cost of ownership? Do the government tax credits make a difference? (4 pt) (0) Electricity prices in other states: For comparison, look at the most recent data on electricity prices from the U.S. Energy Information Administration table graphe Find the column with the most recent (May 2018) data for residential electricity customers. Which state has the highest electricity price? Which state has the lowest? What are the electricity prices in these two states? (4 pt) (k) Comparing costs across state: How much would it cost to fuel the electric-power Leaf if your aunt lived in the state with cheapest electricity instead of Arizona? Does that change your answer to (d)? (6 pt)
Answer #1

Answer:

F)

Car Fuel Efficiency B Total Fuel Used=12000/B
Toyota Yaris 33 miles/ Gallon 12000/33=363.63 gallons
Toyota Prius 50 miles /gallon 12000/50=240 gallons
Nissan Leaf 3 miles /Kwh 12000/3=4000 kwh

G)

Car Total Fuel Used Fuel Cost
Toyota Yaris 363.63 gallons 2*363.63=$727.26
Toyota Prius 240 gallons 2*240=$480
Nissan Leaf 4000 kwh 13*4000/100=$520

From Above table it is clear that Toyata Prius is the cheapest to Drive.

H)

Total Cost for ownership= Purchase cost+5* Fuel cost-Residual Value-Tax Credit

Car Fuel Cost Total Ownership Cost
Toyota Yaris 2*363.63=$727.26 15000+5*727.26-10000=$8,636.3
Toyota Prius 2*240=$480 24000+5*480-2500-12000=$11,900
Nissan Leaf 13*4000/100=$520 30000+5*520-6000-7500=$19,100

I)

From above calculation it is clear that Toyota Yaris is cheapest car and here tax credit is zero for this car. So Tax credit does not make any difference.